OpenStudy (anonymous):

The compound interest on a certain sum of money at 5% p.a. for 2 years is Rs. 246 Calculate the Simple Interest on the same sum for three years.

OpenStudy (amistre64):

5% pa, what does the pa represent?

OpenStudy (amistre64):

246 = P(1+.05/n)^(2n) 246(1+.05/n)^(-2n) = P simple interest = 246(1+.05/n)^(-2n) * .05 * 3 maybe?

OpenStudy (amistre64):

ahh, pa, per annum (once a year so n=1)

OpenStudy (anonymous):

i'm asking u n u r asking me instd? :p r'nt u confident?

OpenStudy (amistre64):

this is a site for studying, not for direct answers. Therefore its best if you can spot any errors in assumptions that I am making ....

OpenStudy (amistre64):

the vernacular tends to throw me at times .... and it helps that you have more context related to a solution than I do

OpenStudy (amistre64):

since pa relates to a compounding of once a year, then the "n" variable is just 1 simple interest is a simple formula: I = Prt using the calculated P, and the stated r and t ... i cant see anywhere that I might have missed it